OpenStudy (anonymous):

With the following information given: y=x^2 , how do i find the coordinates of the vertex of this PARABOLA?

OpenStudy (anonymous):

x^2 cannot be negative - minimum value = 0 at (0,0) vertex is at (0,0)

OpenStudy (anonymous):

how did you find out the vertex coordinates???

OpenStudy (anonymous):

can you take step by step please.

OpenStudy (anonymous):

\[y=4x^2\]

OpenStudy (anonymous):

you can use that example ^^

OpenStudy (anonymous):

the vertex will be at the minimum or maximum value of y in this case minimum value is 0 as a square cannot have a real neagative value when y = 0 x = 0 therefore the vertex is at (0,0)

OpenStudy (anonymous):

hmm use y=4x^2 this time

OpenStudy (anonymous):

y = 4x^2 same argument holds for this - minmumm value of 4x^2 = 0 and when y = 0 x= 0 vertex is at (0,0)

OpenStudy (anonymous):

what do you mean minimum value of 4x^2 = 0 i don’t understand that!!

OpenStudy (anonymous):

well a square value ie in this case x^2 cannot be negative eg -2 & -2 = +4 3 * 3 = 9 so the lowest value a square can be is 0 0 * 0 = 0 and anything times 0 eg 4 * 0 = 0 therefore lowestt value of y^2, 5x^2 ax^2 etc is 0

OpenStudy (anonymous):

ok well what if it’s y=x^2 + 12

OpenStudy (anonymous):

in this case the vertex is at y = 0 + 12 = 12 ie at (0,12) for the general form ax^2 + bx + c thex coordinate of the vertex is - b / 2a

OpenStudy (anonymous):

i haven’t taken that equation yet i only know y=a(x-h) +k

OpenStudy (anonymous):

eg for x^2 + 3x - 6 the vertex is at x = -3/2*1 = -3/2 when x = -3/2 y = (-3/2)^2 + 3 (-3/2) - 6 = -9/4 - 9/2 - 6 = -51/4 so vertex is at (-3/2, -51/4)

OpenStudy (anonymous):

but idk that equation

OpenStudy (anonymous):

i was only taught for now the y = a(x-h)^2 +k

OpenStudy (anonymous):

oh ok this the the vertex form of a quadratic when x = h the a(x-h)^2 = 0 and then y = k

OpenStudy (anonymous):

ok so whenever i see any quadratic equation i grab the H value and the K value and then i WHATEVER THE H AND K VALUES ARE, THATS THE VERTEX COORDINATES?

OpenStudy (anonymous):

and in this form the vertex is (h,k)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

oh but its -h, do i ignore the negative sign on the h?

OpenStudy (anonymous):

say i have y=5(x-3)^2 , the vertex coordinates are (3,0) OR (-3,0)

OpenStudy (anonymous):

in this case the vertex is at (3,0) if y = 5(x+3)^2 the vertex is at (-3,0) because the -3 makes (x+3) = 0

OpenStudy (anonymous):

ohhhhhh i sees now bro.

OpenStudy (anonymous):

so just to get this clear:

OpenStudy (anonymous):

i’m gonna write out a conclusion and i want you to tell me if it’s correct. ok?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

To find out the coordinates of a vertex algebraically, you must take the value of h and replace it in for X, and take k and replace it in for Y. if h is positive, the number will be negative. if h is negative, the value will be positive. If there is no H, the vertex point x=0 . if there is no k, y=0 in the vertex.

OpenStudy (anonymous):

example: y=3(x-3)^2 + 17 vertex point: (3,17) <<<<--- am i right??

OpenStudy (anonymous):

absolutely right and for y=3(x+3)^2 - 17 what is the vertex?

OpenStudy (anonymous):

(3,-17) ?

OpenStudy (anonymous):

nearly : -17 is right but wahat about the x coordinate?

OpenStudy (anonymous):

ohhhh yeah its -3 i forgot yup

OpenStudy (anonymous):

cause h is naturally negative

OpenStudy (anonymous):

so 2 negatives form a positive.

OpenStudy (anonymous):

yup or another way to look at it is (x+3) must be = 0 therefore x must be -3

OpenStudy (anonymous):

ohhh thats a good way too. anyways, thank you so much for your help. if i could give you 100 medals, i would. i really appreciate it, i now understand it because of you! thanks :D

OpenStudy (anonymous):

thats ok

OpenStudy (anonymous):

set first derivative in terms of x equal to zero and solve for x. dx/d= 2x for 2x=0, x=0, plug in 0 for the original equation and y also equals zero, answer: (0,0)

OpenStudy (anonymous):

ty